# Modulo arithmetic of function spaces: Subset hyperspaces as quotients of function spaces

### v1

@@ 4. Concrete quotient-realization and metrization of compact-subset hyperspaces
(i) Pick any open neighborhood $O\in\mathcal{U}_{\lambda}$ of $f(Z)$. Consider the system of open sets $\mathcal{O}:=\{O_{y}=O\ni f(z)\}_{z\in Z}=\{O\}$. Then there exists $\alpha^{\mathcal{O}}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O$, which implies $\bigcup_{\alpha\geq\alpha^{\mathcal{O}}}f_{\alpha}(Z)\subset O$. (ii) Let $O\in\mathcal{U}_{\lambda}$ be an open set such that $O\cap f(Z)\neq\emptyset$ (i.e., $f(Z)\in O^{-}$). Then some $f(z)\in O$. Since $f_{\alpha}(z)\rightarrow f(z)$, there is $\alpha_{z}^{O}$ such that $\{f_{\alpha}(z)\}_{\alpha\geq\alpha_{z}^{O}}\subset O$, i.e., $f_{\alpha}(Z)\cap O\neq\emptyset$ for all $\alpha\geq\alpha_{z}^{O}$. (iii) By Note 1 of Definition 4.1, $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ implies $f_{\alpha}|_{Z}\stackrel{{\scriptstyle\tau_{p}}}{{\longrightarrow}}f|_{Z}$. So, the conclusion follows by (i) and (ii). ∎
@@
Let $X=(X,\mathcal{U})$ be a chained-space, $Y$ a space, $\mathcal{F}\subset X^{Y}$ a $q$-full finitely $q$-stable subspace consisting of closed image maps, and $\mathcal{Z}\subset\mathcal{P}^{\ast}(Y)$ a cover of $Y$ (i.e., $\bigcup\mathcal{Z}=Y$), such that (i) $\mathcal{F}$ is $\tau_{uc}$-closed, (ii) $(\mathcal{F},\tau_{uc|_{\mathcal{Z}}})$ is an rc-space, and (iii) every $\tau_{uc|_{\mathcal{Z}}}$-convergent cr-net in $\mathcal{F}$ has a $\tau_{uc}$-convergent subnet (with convergence limit in $X^{Y}$ in general).
Then the topology $\tau_{\pi}:=\tau_{uc|_{\mathcal{Z}}}$ on $\mathcal{F}\subset X^{Y}$ is a $\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle$-compatible swrc-topology. Moreover, $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{v})$ is continuous (i.e., $\tau_{v}\subset\tau_{\pi q}$ in $Cl_{Y}(X)$ ).
NB: When $Y$ is finite, it suffices (by Theorem 3.17) to replace $\tau_{uc|_{\mathcal{Z}}}$ with $\tau_{p}$, in which case, Tychonoff’s product theorem ensures that the requirements (i) and (ii) are no longer needed.
@@
We need to verify the necessary requirement for $\tau_{\pi}$ in Definition 3.7. By Note 3 of Definition 4.1 and the equality $\bigcup\mathcal{Z}=Y$, we have $\tau_{p}\subset\tau_{uc|_{\mathcal{Z}}}$. Since $\mathcal{F}$ consists of closed image maps (so $q(f)=f(Y)$), we also have $q^{-1}(\tau_{v})\subset\tau_{uc|_{\mathcal{Z}}}$ by Lemma 4.3(iii).
@@
    $\tau_{\pi}$ is an rc-topology (hence a wrc-topology) on $\mathcal{F}$: To see this, let $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ be a cr-net. Since $(\mathcal{F},\tau_{uc|_{\mathcal{Z}}})$ is an rc-space and $\mathcal{F}$ is $\tau_{uc}$-closed, some subnet $f_{\alpha(\beta)}\stackrel{{\scriptstyle\tau_{uc}}}{{\longrightarrow}}f\in\mathcal{F}$. Hence $f_{\alpha(\beta)}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}}}}}{{\longrightarrow}}f\in\mathcal{F}$, since $\tau_{uc|_{\mathcal{Z}}}\subset\tau_{uc}$.
    NB: It is clear that if $Y$ is finite, in which case $\tau_{uc|_{\mathcal{Z}}}\stackrel{{\scriptstyle(\ast)}}{{=}}\tau_{p}\stackrel{{\scriptstyle(\ast)}}{{=}}\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle=\langle\tau_{p}\cup q^{-1}(\tau_{v}^{+})\rangle$ (where the equalities ($\ast$) hold because pointwise convergence automatically implies uniform convergence, which in turn implies $q^{-1}(\tau_{v})$-convergence by Lemma 4.3(iii)), then the conclusion no longer requires the hypotheses (ii) and (iii).
@@
The existence of a swrc-topology on a $q$-full $\mathcal{F}\subset X^{Y}$ (say as in Theorem 4.4) allows us to concretely establish metrization (in Theorems 4.8,4.12) of indexed compact-subset hyperspaces $K_{Y}(X)\subset K(X)$ of a metrizable space $X$.
@@
Let $X$ be a space and $Y$ a set. If $\mathcal{A}\subset(K_{Y}(X),\tau_{v})$ is compact, then so is $K:=\bigcup_{A\in\mathcal{A}}A\subset X$. Hence, if $\mathcal{B}\subset(K_{Y}(X),\tau_{v})$ is contained in a compact subset of $(K_{Y}(X),\tau_{v})$ then $L:=\bigcup_{B\in\mathcal{B}}B$ is contained in a compact subset of $X$.
@@
[…]
@@
Let $X=(X,d)$ be a metric space and $Y$ a set. If $\mathcal{C}\subset\big{(}K_{Y}(X),d_{H}\big{)}$ is compact then $K:=\bigcup_{C\in\mathcal{C}}C\subset X$ is compact.
We review the proof of the following well known result ([IllanNadl1999, Theorem 3.1]).
@@
If $X$ is a metrizable space, then so is $(K(X),\tau_{v})$. Moreover, if $X\cong(X,d)$, then $(K(X),\tau_{v})\cong(K(X),d_{H})$. Conversely, if $X$ is $T_{1}$ (i.e., all points of $X$ are closed) and $K(X)$ is metrizable, then $X$ is metrizable (as a subspace of a metrizable space).
@@
Let $X=(X,d)$. (i) Showing $\tau_{v}\subset\tau_{d_{H}}$ in $K(X)$: Let $O\subset X$ be open. We need to show $O^{+}$ and $O^{-}$ are in $\tau_{d_{H}}$. First, let $A_{n}\in(O^{+})^{c}=\{K\in K(X):K\cap O^{c}\neq\emptyset\}=(O^{c})^{-}$ such that $A_{n}\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}A$. Then with $a_{n}\in A_{n}\cap O^{c}\subset A\cup\bigcup_{n}A_{n}$ (a compact set), let $d(a_{f(n)},c)\rightarrow 0$, where $c\in O^{c}$ (a closed set). Then
$\operatorname{dist}(c,A)$ $=\inf_{a\in A}d(c,a)\leq\inf_{a\in A}[d(c,a_{f(n)})+d(a_{f(n)},a)]=d(c,a_{f(n)})+\operatorname{dist}(a_{f(n)},A)$
$\leq d(c,a_{f(n)})+d_{H}(A_{f(n)},A)\rightarrow 0,$
$~{}~{}\Rightarrow~{}~{}c\in A~{}~{}\Rightarrow~{}~{}O^{c}\cap A\neq\emptyset,~{}~{}\Rightarrow~{}~{}A\in(O^{+})^{c},$
@@
Second, let $A_{n}\in(O^{-})^{c}=\{K\in K(X):K\subset O^{c}\}=(O^{c})^{+}$ such that $A_{n}\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}A$. Fix $a\in A$. Then, for some $a_{n}=a_{n}(a)\in A_{n}$, we have $d(a,a_{n})=\operatorname{dist}(a,A_{n})\leq d_{H}(A,A_{n})\rightarrow 0$. Since $a_{n}\in A\cup\bigcup_{n}A_{n}$ (a compact set), let $d(a_{n},c)\rightarrow 0$ for some $c\in O^{c}$ (since $A_{n}\subset O^{c}$ and $O^{c}$ is a closed set). Then
$d(a,c)\leq d(a,a_{n})+d(a_{n},c)\rightarrow 0,~{}~{}\Rightarrow~{}~{}a=c,~{}~{}\Rightarrow~{}~{}A\subset O^{c},$
@@
$\textstyle d_{H}(A,A_{\alpha})$ $=\max_{A\leftrightarrow A_{\alpha}}\max_{a\in A}\operatorname{dist}(a,A_{\alpha})\leq\max_{A\leftrightarrow A_{\alpha}}\max_{a\in A}\operatorname{dist}(a,A_{\alpha}\cap O_{i_{a}})|_{a\in O_{i_{a}}}\leq\varepsilon^{\mathcal{O}}.$
@@
If $X=(X,d)$ is a metric space and $Y$ a finite set, then  ${}(Cl_{Y}(X),\tau_{pq})=(Cl_{Y}(X),d_{H})$.
@@
Since $Cl_{Y}(X)=FS_{|Y|}(X)\subset K(X)$, it follows from Lemma 4.7 that “$(Cl_{Y}(X),\tau_{v})=(Cl_{Y}(X),d_{H})$”, where by Theorem 3.17, $\tau_{pq}=\tau_{v}$ in $Cl_{Y}(X)$. ∎
@@
Let $X$ be a set, $(Y,\tau)$ a space, $f:X\rightarrow Y$ a map, and $y_{\alpha}\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}y$ a convergent net in $Y$. If $x_{\alpha}\in f^{-1}(y_{\alpha})$ and $x\in f^{-1}(y)$, then  $x_{\alpha}\stackrel{{\scriptstyle f^{-1}(\tau)}}{{\longrightarrow}}x$ in $X$ (where $f^{-1}(\tau):=\{f^{-1}(U):U\in\tau\}$). (footnote111111 Since $f:(X,f^{-1}(\tau))\rightarrow(Y,\tau)$ is continuous, for every convergent net $x_{\alpha}\stackrel{{\scriptstyle f^{-1}(\tau)}}{{\longrightarrow}}x$ in $X$, we have $f(x_{\alpha})\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}f(x)$ in $Y$. ).
@@
Let $\mathcal{C}\subset K_{Y}(X)$ be a $\tau_{\pi q}$-closed set. Consider a sequence $\left\{q(f_{n})\right\}_{n}\subset\mathcal{C}$ such that $q(f_{n})\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}q(f)$ for some $q(f)\in BCl_{Y}(X)$. We want to show that $q(f)\in\mathcal{C}$ (in which case $\mathcal{C}$ is also $d_{H}$-closed). In $\mathcal{F}$, let $g_{n}\in q^{-1}\left(q(f_{n})\right)\subset q^{-1}(\mathcal{C})$, i.e., $q(g_{n})=q(f_{n})$ (footnote [^12] A natural choice here is $g_{n}:=f_{n}$. So, when $Y$ is finite, it is enough to take $\tau_{\pi}:=\tau_{p}$ and thereby automatically get a pointwise convergent subnet of the $cr$-sequence $\{g_{n}\}$ by Tychonoff’s product theorem.).
Since $\left\{q(f)\right\}\cup\left\{q(f_{n})\right\}_{n}\subset\big{(}K_{Y}(X),d_{H}\big{)}$ is compact, $K=q(f)\cup\bigcup_{n}q(f_{n})\subset(X,d)$ is compact by Lemma 4.6. So, by the definition of $\tau_{\pi}$, the cr-sequence $\{g_{n}:Y\rightarrow q(f_{n})\subset K\subset(X,d)\}_{n}$ has a $\tau_{\pi}$-convergent subnet $\{g_{n(\alpha)}\}\stackrel{{\scriptstyle\tau_{\pi}}}{{\longrightarrow}}h$, where $h\in q^{-1}(\mathcal{C})$ since $q^{-1}(\mathcal{C})$ is $\tau_{\pi}$-closed. By the continuity of (4.1),

### v2

@@ 4. Concrete quotient-realization and metrization of compact-subset hyperspaces
(i) Pick any open neighborhood $O\in\mathcal{U}_{\lambda}$ of $f(Z)$. Consider the system of open sets $\mathcal{O}:=\{O_{y}=O\ni f(z)\}_{z\in Z}=\{O\}$. Then there exists $\alpha^{\mathcal{O}}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O$, which implies $\bigcup_{\alpha\geq\alpha^{\mathcal{O}}}f_{\alpha}(Z)\subset O$. (ii) Let $O\in\mathcal{U}_{\lambda}$ be an open set such that $O\cap f(Z)\neq\emptyset$ (i.e., $f(Z)\in O^{-}$). Then some $f(z)\in O$. Since $f_{\alpha}(z)\rightarrow f(z)$, there is $\alpha_{z}^{O}$ such that $\{f_{\alpha}(z)\}_{\alpha\geq\alpha_{z}^{O}}\subset O$, i.e., $f_{\alpha}(Z)\cap O\neq\emptyset$ for all $\alpha\geq\alpha_{z}^{O}$. (iii) By Note 2 of Definition 12, $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ implies $f_{\alpha}|_{Z}\stackrel{{\scriptstyle\tau_{p}}}{{\longrightarrow}}f|_{Z}$. So, the conclusion follows by (i) and (ii). ∎
@@
Let $X=(X,\mathcal{U})$ be a chained-space, $Y$ a space, and $\mathcal{Z}\subset\mathcal{P}^{\ast}(Y)$ a cover of $Y$ (i.e., $\bigcup\mathcal{Z}=Y$). Let $\mathcal{F}\subset X^{Y}$ be a $q$-full (and finitely $q$-stable) subspace such that (i) $\overline{f(Z)}=f(Z)$ for all $\{f\in\mathcal{F},Z\in\mathcal{Z}\}$, and (ii) $(\mathcal{F},\tau)$ is an rc-space with $\tau_{uc|_{\mathcal{Z}}}\subset\tau$ (for example, we could take $\tau:=\tau_{uc}$). Then the topology $\tau_{\pi}:=\tau_{uc|_{\mathcal{Z}}}$ on $\mathcal{F}\subset X^{Y}$ is a $\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle$-compatible wrc-topology and, moreover, $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{v})$ is continuous (i.e., $\tau_{v}\subset\tau_{\pi q}$ in $Cl_{Y}(X)$ ).
NB: When $Y$ is finite, it suffices (by Theorem 3.3) to replace $\tau_{uc|_{\mathcal{Z}}}$ with $\tau_{p}$, in which case, Tychonoff’s product theorem ensures that (i) and (ii) are no longer needed.
@@
We need to verify the necessary requirement for $\tau_{\pi}$ in Definition 7. By Note 4 of Definition 12 and the equality $\bigcup\mathcal{Z}=Y$, we have $\tau_{p}\subset\tau_{uc|_{\mathcal{Z}}}$. Since $\mathcal{F}$ satisfies (i), we also have $q^{-1}(\tau_{v})\subset\tau_{uc|_{\mathcal{Z}}}$ by Lemma 4.1(iii). Therefore $\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle\subset\tau_{uc|_{\mathcal{Z}}}$.
@@
    $\tau_{\pi}$ is an rc-topology (hence a wrc-topology) on $\mathcal{F}$: Indeed, for any cr-net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$, the rc-space $(\mathcal{F},\tau)$ gives a subnet $f_{\alpha(\beta)}\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}f\in\mathcal{F}$. So $f_{\alpha(\beta)}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}}}}}{{\longrightarrow}}f\in\mathcal{F}$ (since $\tau_{uc|_{\mathcal{Z}}}\subset\tau$).
    NB: It is clear that if $Y$ is finite, in which case $\tau_{uc|_{\mathcal{Z}}}\stackrel{{\scriptstyle(\ast)}}{{=}}\tau_{p}\stackrel{{\scriptstyle(\ast)}}{{=}}\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle=\langle\tau_{p}\cup q^{-1}(\tau_{v}^{+})\rangle$ (where the equalities ($\ast$) hold because pointwise convergence automatically implies uniform convergence, which in turn implies $q^{-1}(\tau_{v})$-convergence by Lemma 4.1(iii)), then the conclusion no longer requires (i) and (ii).
@@
Metrization of compact-subset hyperspaces
The existence of a swrc-topology on a $q$-full $\mathcal{F}\subset X^{Y}$ (say as in Theorem 4.2) allows us to concretely establish metrizability (in Theorems 4.6 and 4.9) of indexed compact-subset hyperspaces $K_{Y}(X)\subset K(X)$ of a metrizable space $X$.
@@
Let $X$ be a space and $Y$ a set. If $\mathcal{A}\subset(K_{Y}(X),\tau_{v})$ is compact, then so is $K:=\bigcup_{A\in\mathcal{A}}A\subset X$. Hence, if $\mathcal{B}\subset(K_{Y}(X),\tau_{v})$ is contained in a compact subset of $(K_{Y}(X),\tau_{v})$, then $L:=\bigcup_{B\in\mathcal{B}}B$ is contained in a compact subset of $X$.
@@
Consider a net $x_{\alpha}\in K$, and let $x_{\alpha}\in A_{\alpha}\in\mathcal{A}$. By the compactness of $\mathcal{A}$, let a subnet $A_{\alpha(\beta)}\stackrel{{\scriptstyle\tau_{v}}}{{\longrightarrow}}A_{0}\in\mathcal{A}$ (i.e., for any $\mathcal{O}:=[\{O_{1},...,O_{n}\}]_{v}\ni A_{0}$, some tail $\{A_{\alpha(\beta)\geq\alpha(\beta^{\mathcal{O}})}\}\subset[\{O_{1},...,O_{n}\}]_{v}$). We will show there is $x\in A_{0}$ such that $x_{\alpha(\beta)}\rightarrow x$. On the contrary, suppose that for all $x\in A_{0}$, $x_{\alpha(\beta)}\not\!\!\longrightarrow x$, i.e., there exists an open set $O_{x}\ni x$ and a subnet $\{x_{\alpha\circ\beta_{x}(\gamma)}\}$ that avoids $O_{x}$. In particular, since $n$ is finite, (a tail of) some subnet $\{x_{\alpha\circ\beta(\gamma)}\}$ avoids $O_{1},...,O_{n}$. But $\{A_{\alpha\circ\beta(\gamma)\geq\alpha\circ\beta(\gamma^{\mathcal{O}})}\}\subset[\{O_{1},...,O_{n}\}]_{v}$ implies $\{x_{\alpha\circ\beta(\gamma)\geq\alpha\circ\beta(\gamma^{\mathcal{O}})}\}\subset\bigcup_{i}O_{i}$ (a contradiction). (footnote [^9] Alternative proof: Let $\{O_{i}:i\in I\}$ be an open cover of $K$ (hence also an open cover of each $A\in\mathcal{A}$) in $X$. For each $A\in\mathcal{A}$, let $\{O^{A}_{i_{j}}:j\in I_{A}\}\subset\{O_{i}:i\in I\}$ be a finite subcover of AA (where wlog $A\cap O_{i_{j}}^{A}\neq\emptyset$ for all $j$). Then $A\in\langle O^{A}_{i_{j}}:j\in I_{A}\rangle_{v}$, and so $\big\{[\{O^{A}_{i_{j}}:j\in I_{A}\}]_{v}\big\}_{A\in\mathcal{A}}$ is a $\tau_{v}$-open cover of $\mathcal{A}$, which therefore has a finite $\tau_{v}$-subscover $\big\{[\{O^{A}_{i_{j}}:j\in I_{A}\}]_{v}\big\}_{A\in F}$, for some finite set $F\subset\mathcal{A}$. Hence $\{O^{A}_{i_{j}}:j\in I_{A},A\in F\}$ is a finite subcover of $K$ in $X$.). ∎
@@
Let $X=(X,d)$ be a metric space and $Y$ a set. If $\mathcal{C}\subset\big(K_{Y}(X),d_{H}\big)$ is compact, then $K:=\bigcup_{C\in\mathcal{C}}C\subset X$ is compact.
We review the proof of the following well known result.
@@
If $X$ is a metrizable space, then so is $(K(X),\tau_{v})$. Moreover, if $X\cong(X,d)$, then $(K(X),\tau_{v})\cong(K(X),d_{H})$. Conversely, if $X$ is $T_{1}$ (i.e., all singletons of $X$ are closed) and $K(X)$ is metrizable, then $X$ is metrizable (as a subspace of a metrizable space).
@@
Let $X=(X,d)$. (i) Showing $\tau_{v}\subset\tau_{d_{H}}$ in $K(X)$: Let $O\subset X$ be open. We need to show $O^{+}$ and $O^{-}$ are in $\tau_{d_{H}}$. First, let $A_{n}\in(O^{+})^{c}=\{K\in K(X):K\cap O^{c}\neq\emptyset\}=(O^{c})^{-}$ such that $A_{n}\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}A$. Then with $a_{n}\in A_{n}\cap O^{c}\subset A\cup\bigcup_{n}A_{n}$ (a compact set), let $d(a_{f(n)},e)\rightarrow 0$, where $e\in O^{c}$ (a closed set). Then
$d(e,A)$ $=\inf_{a\in A}d(e,a)\leq\inf_{a\in A}[d(e,a_{f(n)})+d(a_{f(n)},a)]=d(e,a_{f(n)})+d(a_{f(n)},A)$
$\leq d(e,a_{f(n)})+d_{H}(A_{f(n)},A)\rightarrow 0,$
$~~\Rightarrow~~e\in A~~\Rightarrow~~O^{c}\cap A\neq\emptyset,~~\Rightarrow~~A\in(O^{+})^{c},$
@@
Second, let $A_{n}\in(O^{-})^{c}=\{K\in K(X):K\subset O^{c}\}=(O^{c})^{+}$ such that $A_{n}\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}A$. Fix $a\in A$. Then, for some $a_{n}=a_{n}(a)\in A_{n}$, we have $d(a,a_{n})=d(a,A_{n})\leq d_{H}(A,A_{n})\rightarrow 0$. Since $a_{n}\in A\cup\bigcup_{n}A_{n}$ (a compact set), let $d(a_{n},e)\rightarrow 0$ for some $e\in O^{c}$ (since $A_{n}\subset O^{c}$ and $O^{c}$ is a closed set). Then
$d(a,e)\leq d(a,a_{n})+d(a_{n},e)\rightarrow 0,~~\Rightarrow~~a=e,~~\Rightarrow~~A\subset O^{c},$
@@
$\textstyle d_{H}(A,A_{\alpha})$ $=\max_{A\leftrightarrow A_{\alpha}}\max_{a\in A}d(a,A_{\alpha})\leq\max_{A\leftrightarrow A_{\alpha}}\max_{a\in A}d(a,A_{\alpha}\cap O_{i_{a}})|_{a\in O_{i_{a}}}\leq\varepsilon^{\mathcal{O}}.$
@@
If $X=(X,d)$ is a metric space and $Y$ a finite set, then  $(Cl_{Y}(X),\tau_{pq})=(Cl_{Y}(X),\tau_{v})=(Cl_{Y}(X),d_{H})$.
@@
Since $Cl_{Y}(X)=FS_{|Y|}(X)\subset K(X)$, it follows from Lemma 4.5 that $(Cl_{Y}(X),\tau_{v})=(Cl_{Y}(X),d_{H})$”, where by Theorem 3.3, $\tau_{pq}=\tau_{v}$ in $Cl_{Y}(X)$. ∎
@@
Let $X$ be a set, $(Y,\tau)$ a space, $f:X\rightarrow Y$ a map, and $y_{\alpha}\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}y$ a convergent net in $Y$. If $x_{\alpha}\in f^{-1}(y_{\alpha})$ and $x\in f^{-1}(y)$, then  $x_{\alpha}\stackrel{{\scriptstyle f^{-1}(\tau)}}{{\longrightarrow}}x$ in $X$ (where $f^{-1}(\tau):=\{f^{-1}(U):U\in\tau\}$). (footnote [^10] Since $f:(X,f^{-1}(\tau))\rightarrow(Y,\tau)$ is continuous, for every convergent net $x_{\alpha}\stackrel{{\scriptstyle f^{-1}(\tau)}}{{\longrightarrow}}x$ in $X$, we have $f(x_{\alpha})\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}f(x)$ in $Y$. ).
@@
Let $\mathcal{C}\subset K_{Y}(X)$ be a $\tau_{\pi q}$-closed set. Consider a sequence $\left\{q(f_{n})\right\}_{n}\subset\mathcal{C}$ such that $q(f_{n})\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}q(f)$ for some $q(f)\in BCl_{Y}(X)$. We want to show that $q(f)\in\mathcal{C}$ (in which case $\mathcal{C}$ is also $d_{H}$-closed). In $\mathcal{F}$, let $g_{n}\in q^{-1}\left(q(f_{n})\right)\subset q^{-1}(\mathcal{C})$, i.e., $q(g_{n})=q(f_{n})$ (footnote [^11] A natural choice here is $g_{n}:=f_{n}$. So, when $Y$ is finite, it is enough to take $\tau_{\pi}:=\tau_{p}$ and thereby automatically get a pointwise convergent subnet of the crcr-sequence $\{g_{n}\}$ by Tychonoff’s product theorem (thereby obtaining an alternative proof of Theorem 4.6 without Theorem 3.3).).
Since $\left\{q(f)\right\}\cup\left\{q(f_{n})\right\}_{n}\subset\big(K_{Y}(X),d_{H}\big)$ is compact, $K=q(f)\cup\bigcup_{n}q(f_{n})\subset(X,d)$ is compact by Lemma 4.4. So, by the definition of $\tau_{\pi}$, the cr-sequence $\{g_{n}:Y\rightarrow q(f_{n})\subset K\subset(X,d)\}_{n}$ has a $\tau_{\pi}$-convergent subnet $\{g_{n(\alpha)}\}\stackrel{{\scriptstyle\tau_{\pi}}}{{\longrightarrow}}h$, where $h\in q^{-1}(\mathcal{C})$ since $q^{-1}(\mathcal{C})$ is $\tau_{\pi}$-closed. By the continuity of (4),